A worker lift up of 3.5kg to height of 1.80m each 3seconds find the work done in one hour
1
Expert's answer
2019-05-27T11:35:28-0400

The work done per one second

"P=\frac{W}{t}=\frac{mgh}{t}"

"=\frac{3.5\:\rm{kg}\times 10\:\rm{m/s^2}\times 1.80\:\rm{m}}{3\:\rm{s}}=21\:\rm{W}"

The work done in one hour

"W=Pt=21\:\rm{W}\times 3600\:\rm{s}=75600\:\rm{J}=75.6\:\rm{kJ}"